OpenStudy (anonymous):

Find the surface area of a rectangular prism that is 16 inches long, 12 inches wide, and 5 inches high. A. 960 in.2 B. 689 in.2 C. 714 in.2 D. 664 in.2

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

@ranga

jigglypuff314 (jigglypuff314):

so... (2*5*12)+(2*5*16)+(2*12*16)

OpenStudy (anonymous):

714 in 2

jigglypuff314 (jigglypuff314):

that's not what I got... 2*5*12=120 2*5*16=160 2*12*16=384 then 120+160+384=

OpenStudy (anonymous):

664 in 2

jigglypuff314 (jigglypuff314):

yes

OpenStudy (anonymous):

ok thanks, my last one The length of a rectangular room is 7.9 m and its width is 8.6 m. Find the area of the room. A. 73.96 m2 B. 62.41 m2 C. 67.94 m2 D. 33 m2

jigglypuff314 (jigglypuff314):

7.9*8.6=

OpenStudy (anonymous):

67.94 m 2

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

thank you :]